# Validation tests of Gaussian boson samplers with photon-number resolving detectors

### v1

@@ V.2 Grouped count probability comparisons
First, we compare the theoretical one-dimensional GCPs of the ideal ground truth distribution to the experimental total counts. Similar comparisons were performed in [5], however it was found that at least for the $288$-mode data set the ideal ground truth total count distribution was incorrect [48], leading to inaccurate comparisons. Results of phase-space simulations are presented in Fig.(2a) for the $216$-mode HS data set and Fig.(3a) for the $288$-mode data set. In both cases, the simulated GCP ideal ground truth distributions appear to closely resemble the experimental data. However, upon closer inspection there are significant differences between the two distributions.
@@
This is first observed numerically from statistical testing, where the $216$-mode HS experiment produces a chi-square result of $\chi_{EI}^{2}/k\approx 54$ with $Z_{EI}\approx 70$ and count patterns from the $288$-mode experiment output $\chi_{EI}^{2}/k\approx 411$ with $Z_{EI}\approx 167$. Although both experimental GCPs are far from their pure squeezed state distributions, the $288$-mode distribution is $98\sigma_{\mathcal{N}}$ further from its expected normally distributed mean. These numerical differences are also reflected graphically by plotting the normalized difference between distributions $\Delta\mathcal{G}_{S}^{(M)}(m)/\sigma_{i}=\left(\mathcal{G}_{E,i}-\bar{\mathcal{G}}_{I,i}\right)/\sigma_{i}$, which is plotted in Fig.(2b) for the $216$-mode HS data set and Fig.(3d) for the $288$-mode data set.
@@
Figure 2: Comparisons of one-dimensional GCPs $\mathcal{G}_{\{1,\dots,M\}}^{(M)}(m)$ versus grouped counts $m$ for experimental data obtained from the $216$-mode HS Borealis experiment. a) Data is compared to positive-P simulations with pure squeezed state inputs, where simulations are performed for an ensemble size of $E_{S}=1.2\times 10^{6}$, and the corresponding normalized difference between distributions is plotted in b). In c) experimental patterns are now compared to the simulated GCP using thermalized squeezed state inputs with $\epsilon=0.0510\pm 0.0005$. A transmission matrix measurement correction is also applied as $t\boldsymbol{T}$, where $t=0.9941\pm 0.0005$, to improve the fit. d) The normalized difference between thermalized and experimental GCPs, which display noticeable improvement when compared to pure squeezed state simulations.
@@
Once a thermalization component $\epsilon$ is added to the input states and a transmission matrix measurement correction $t$ is applied as $t\boldsymbol{T}$, the positive-P simulated GCPs agree within sampling error to the experimental distributions. The $216$-mode HS data set requires $\epsilon=0.0510\pm 0.0005$ and $t=0.9941\pm 0.0005$ to output $\chi_{ET}^{2}/k\approx 0.88\pm 0.02$ and $\left|Z_{ET}\right|\approx 1\pm 0.2$, while the best-fit theoretical GCP for $288$-mode data uses a similar thermalization component $\epsilon=0.0547\pm 0.0005$ and measurement correction $t=0.9848\pm 0.0005$ to produce $\chi_{ET}^{2}/k\approx 0.87\pm 0.02$ and $Z_{ET}\approx\left|1.1\right|\pm 0.2$. This exact agreement between theory and experiment is reflected in the normalized difference plots as seen in Fig.(2d) and Fig.(3d) which have noticeably decreased from their pure squeezed state comparisons.
@@
Despite this excellent agreement between theoretical and experimental GCPs once decoherence and measurement corrections are accounted for, we note that stronger validation tests of quantum advantage based on multidimensional GCPs cannot be performed on these data sets. As the dimension increases, the number of bins generated scales as $\left(Md^{-1}+1\right)^{d}$. If there are too few counts per bin, i.e. the bins become more sparse, the experimental sampling errors of these bins increase dramatically, causing statistical tests to become dominated by sampling errors, outputting artificially small results.
Earlier works [16] on recent threshold detector experiments generating $N_{E}\approx 4\rightarrow 5\times 10^{7}$ count patterns showed these sampling error limitations arose for a GCP dimension of $d=4$, such that reliable validation tests could only be performed for $d<4$. However, the majority of data sets from Borealis generate $N_{E}\approx 1\times 10^{6}$ patterns, causing sampling errors to become dominant at only $d=2$, limiting comparisons to only $d=1$.
The exception to this is the $16$-mode experiment whose $N_{E}\approx 8.4\times 10^{7}$ samples were used benchmark the system architecture through total variation distance tests as the system is small enough that output photo-detection measurements can be replicated exactly. Interestingly, the one-dimensional GCP for this data set is further from its ideal ground truth distribution than most experiments performed using the Borealis quantum network, where $\chi_{EI}^{2}/k\approx 2.5\times 10^{3}$ and $Z_{EI}=163$. Unlike data sets claiming quantum advantage, including decoherence and $\boldsymbol{T}$-matrix corrections is no longer enough to obtain a within sampling error agreement between theoretical and experimental GCPs, instead producing statistical test results of $\chi_{ET}^{2}/k\approx 14\pm 2$ and $Z_{ET}\approx 18\pm 1$.
@@
Figure 3: Description follows Fig.(2) except experimental count patterns are from the $288$-mode Borealis network. Input thermalized squeezed state simulations are performed using a decoherence component of $\epsilon=0.0547\pm 0.0005$ and matrix measurement error $t=0.9848\pm 0.0005$.
@@
Due to the large number of samples, statistical tests of multi-dimensional GCPs can be performed on this data set. A small increase in the GCP dimension to $d=2$ shows the $16$-mode data is even further from both the ideal and non-ideal ground truth distributions once high-order correlations become more statistically prevalent, outputting $Z_{EI}=301$ and $Z_{ET}=86$. An exponential number of comparison tests can be performed by randomly permuting the measured photon count patters [16], where each permutation allows one to compare different binned correlation moments. As some correlation moments may be closer to their theoretical values than others, Z-scores are likely to vary between permutations, as was observed in threshold detector experiment comparisons [16]. Focusing on comparisons with the non-ideal ground truth, a mean increase in the resulting Z-scores is observed from ten permutations, such that $\left\langle Z_{ET}\right\rangle_{\mathcal{RP}}\approx 139$, where $\left\langle\dots\right\rangle_{\mathcal{RP}}$ denotes a mean over random permutations. Although some permutations are closer to the non-ideal ground truth, with the closest being $Z_{ET}\approx 73$, the majority are further, where the largest observed Z-score was $Z_{ET}\approx 235$ from a single permutation.
@@
Although this is not unexpected, such results indicate further systematic errors besides those tested are present. We conjecture that the presence of further systematic errors other than decoherence and measurement errors are also present in the $216$-mode HS and $288$-mode implementations. However the small sample size means such errors cannot be resolved in statistical testing for these data sets, limiting a thorough analysis.
V.3 Photon number moments
@@
However, unlike the GCP comparisons, decoherence and transmission matrix corrections no longer cause the theoretical moments to be within error of the experimental moments, where $Z_{ET}\approx 45$ for the $216$-mode HS data set and the $288$-mode data set is $Z_{ET}\sigma_{\mathcal{N}}\approx 63\sigma_{\mathcal{N}}$ standard deviations away from its expected mean. Similar trends are present for all data sets where, as was the case for GCPs, the $16$-mode moments are furthest from their expected first order moments than any other experiment. We emphasize that statistical testing is more accurate for this data set due to the large number of experimental samples, allowing one to resolve finer differences between theoretical and experimental photon number moments.
@@
Figure 4: a) Input thermalized squeezed state theoretical mean output photon numbers $\left\langle\hat{n}^{\prime}_{j}\right\rangle$ versus mode number $j$ are compared to their experimental counterparts for data from the $216$-mode HS Borealis experiment, where simulations are performed for an ensemble size of $E_{S}=1.2\times 10^{6}$. Decoherence and $\boldsymbol{T}$-matrix parameters are $\epsilon=0.0510\pm 0.0005$ and $t=0.9941\pm 0.0005$ which, although improving the fit with experimental data, no longer cause theory and experiment to agree within sampling error. b) Corresponding normalized difference between marginal moments.
@@
Figure 5: Figure description follows the same as Fig.(4), except data is now from the $288$-mode Borealis experiment where corresponding fitting parameters are $\epsilon=0.0547\pm 0.0005$ and $t=0.9848\pm 0.0005$.
@@ VI Conclusion
This method is computationally efficient, with computational run-times of a few minutes on a 12-core desktop at $\sim 10^{11}$ floating point operations/ sec (Linpack Flops) for one-dimensional GCP simulations of the largest sized networks implemented in the paper. Since the positive P-representation accurately computes correlation moments to any order generated in the photonic network, it is a time and resource efficient method of validation compared to classical algorithms that compute the Hafnian to replicate photo-detection measurements, or tensor network approximations [31]. Increasing the GCP dimension causes an increase in computation time, although this is still significantly faster than other validation methods. Directly calculating $100\times 100$ Hafnians are estimated [31] to take $14$ hours on Fugaku, one of the worlds largest supercomputers in 2024, at $\sim 10^{18}$ Flops. This would compute one probability, but rounding errors are unknown, as the calculation has not occurred. Generating enough classical counts for validation using this approach would take $\sim 10^{16}s$ [31] on Fugaku, indicating an overall speed-up of $>10^{18}$ using the +P method at these network sizes.
@@
We emphasize that our method does not generate classical counts, so it is strictly for validation of the computational task, and does not replicate the required random counts. Validation tests were performed on GBS data from a recent experiment claiming quantum advantage [5]. We showed that within sampling error agreement between theory and experiment for GCPs is obtained once decoherence is included in the input states, and a transmission matrix measurement error correction is applied. There are still large discrepancies in individual channel count probabilities, possibly because these can be calculated very accurately. This result agrees with a trend obtained previously for threshold detector experiments [15, 16].
@@
These comparisons with threshold detector experiments are slightly misleading. Due to the smaller number of experimental samples from data sets claiming quantum advantage using PNR methods, only a small fraction of the possible tests could be performed. This was because experimental sampling errors were too large to apply more sensitive tests. Based on comparisons of the $16$-mode experimental data, whose sample size was large enough to perform accurate statistical testing, we conjecture that more experimental errors are present in the quantum advantage data sets than those tested for in this paper.
@@
In the longer term, due to its computational efficiency and ability to generate any measurable moment, the positive-P based binning method could be used as a means to provide error correction, either by correcting parameter values as shown in this paper, or by giving feedback so they can be adjusted. We suggest that this could provide a more reasonable route to quantum computational advantage than simply using improved hardware, as is also thought to be the case for logic gate based quantum computers.
@@ Acknowledgments
The graphs presented above have results of statistical tests for comparisons of theory and experiment predominantly for the $216$-mode HS and $288$-mode data sets, due to claims of quantum advantage. Here we summarize statistical test results for all data sets in [5], starting with one-dimensional GCPs for the ideal ground truth in Table.(1). We use the subscripts $EI$ and $ET$ to denote comparisons between experimental data and the ideal and non-ideal ground truth distributions, respectively.
A summary of the non-ideal ground truth comparisons are presented in Table.(2), where the $t$ and $\epsilon$ fitting parameters are found using a Nelder-Mead simplex algorithm which minimizes the distance between experimental and theoretical distributions. As stated earlier, experimental data agrees with their theoretical predictions within sampling error for all data sets bar the $16$-mode experiment. Meanwhile comparisons of the mean output photon number moments are presented in Table.(3) for both the ideal and non-ideal ground truth distributions.
@@
Appendix B: Estimating experimental sampling errors of moments
@@
Theoretical sampling errors are estimated numerically by subdividing the total phase-space ensemble as $E_{S}=N_{S}N_{R}$. Breaking the full ensemble of phase-space samples into sub-ensembles is a common numerical procedure for applications of phase-space representations yielding two complementary benefits: efficient multi-core parallel computing, decreasing computation time depending on available resources, and estimates of theoretical sampling errors [16, 41].
The first sub-ensemble corresponds to sampling from the Gaussian phase-space distribution $N_{S}$ times and $N_{R}$ is the number of times the computation is repeated. Provided $N_{R}\gg 1$, the second sub-ensemble allows one to accurately estimate the standard deviation in the ensemble mean as $\sigma_{T,i}=\sigma_{t,i}/\sqrt{N_{R}}$ [47, 50], where $\sigma_{t}$ is the computed sample standard deviation which takes the general form [16, 41]
@@
Figure 6: The exactly computable mean output photon number (solid black line) of $N=50$ thermal states sent into a lossless linear network $\boldsymbol{U}$ is compared to the mean output photon number (solid orange dots) generated from a classical algorithm that replicates PNR photo-detection measurements of classical states using the diagonal P-representation. The $N_{F}=4\times 10^{6}$ fake count patterns are binned to form the observable moments, where the standard deviation of these moments from their exact values $\sigma_{A}$ is the actual sampling error we aim to estimate using Eq.(45).

### v2

@@ V.2 $d=1$ grouped count comparisons
First we compare the theoretical and experimental one-dimensional GCPs of the ideal ground truth distribution. Results of phase-space simulations are presented in Fig.(3a) for the $216$-mode HS data set and Fig.(4a) for the $288$-mode data set. In both cases, the simulated GCP ideal ground truth distributions appear to visually resemble the experimental data.
Visual comparisons of these distributions were also presented in [5]. When comparing our theoretical total count distributions to those obtained in [5], we found that for the $288$-mode data set, the two theoretical distributions differed slightly. After further checks by the original authors [53], the theoretical total count distribution in Fig.(4a) was determined to be the correct ideal ground truth distribution.
@@
Despite the apparent visual agreement between theory and experiment, a quantitative $\chi^{2}$ test shows significant differences between the two distributions. The $216$-mode HS experimental data gives a result of $\chi_{EI}^{2}/k\approx 54$ with $Z_{EI}\approx 70$ Count patterns from the $288$-mode experiment have $\chi_{EI}^{2}/k\approx 411$ with $Z_{EI}\approx 167$. Although both experimental GCPs are far from their pure squeezed state distributions, the $288$-mode distribution is $98\sigma_{\mathcal{N}}$ further from its expected normally distributed mean. These numerical differences are also reflected graphically by plotting the normalized difference between distributions $\Delta\mathcal{G}_{S}^{(M)}(m)/\sigma_{i}=\left(\mathcal{G}_{E,i}-\bar{\mathcal{G}}_{I,i}\right)/\sigma_{i}$, which is plotted in Fig.(3b) for the $216$-mode HS data set and Fig.(4b) for the $288$-mode data set.
@@
Figure 3: Comparisons of one-dimensional GCPs $\mathcal{G}_{\{1,\dots,M\}}^{(M)}(m)$ versus grouped counts $m$ for experimental data obtained from the $216$-mode HS Borealis experiment. a) Data is compared to positive-P simulations with pure squeezed state inputs, where simulations are performed for an ensemble size of $E_{S}=1.2\times 10^{6}$, and the corresponding normalised difference between distributions is plotted in b). In c) experimental patterns are now compared to the simulated GCP using thermalised squeezed state inputs with $\epsilon=0.0510\pm 0.0005$. A transmission matrix measurement correction is also applied as $t\boldsymbol{T}$, where $t=0.9941\pm 0.0005$, to improve the fit. d) The normalised difference between thermalized and experimental GCPs, which display noticeable improvement when compared to pure squeezed state simulations.
@@
Once a thermal component $\epsilon$ is added to the input states and a transmission matrix measurement correction $t$ is applied as $t\boldsymbol{T}$, the positive-P simulated GCPs agree with the experimental distributions within sampling error. The $216$-mode HS data set requires $\epsilon=0.0510\pm 0.0005$ and $t=0.9941\pm 0.0005$ to output $\chi_{ET}^{2}/k\approx 0.88\pm 0.02$ and $\left|Z_{ET}\right|\approx 1\pm 0.2$, while the best-fit theoretical GCP for the $288$-mode data uses a similar thermal fraction of $\epsilon=0.0547\pm 0.0005$ and measurement correction $t=0.9848\pm 0.0005$ to produce $\chi_{ET}^{2}/k\approx 0.87\pm 0.02$ and $Z_{ET}\approx\left|1.1\right|\pm 0.2$. This improved agreement between theory and experiment is reflected in the normalized difference plots as seen in Fig.(3d) and Fig.(4d) which have noticeably decreased from their pure squeezed state comparisons.
@@
Figure 4: Description follows Fig.(3) except experimental count patterns are from the $288$-mode Borealis network. Input thermalised squeezed state simulations are performed using a decoherence component of $\epsilon=0.0547\pm 0.0005$ and matrix measurement error $t=0.9848\pm 0.0005$.
V.3 $d=2$ grouped count comparisons
To determine whether this excellent agreement between theoretical and experimental GCPs once decoherence and measurement corrections are accounted for continues, we increase the binning dimension to $d=2$. In this case, the $216$-mode HS experiment, whose ideal and thermalized $d=2$ distributions are shown in Fig.(5), outputs statistical test results of $Z_{EI}=58$ and $Z_{ET}=9\pm 1$.
Although the $Z_{ET}$ results are $\approx 8\sigma_{\mathcal{N}}$ larger than their $d=1$ counterparts, the ideal ground truth comparisons are $\approx 12\sigma_{\mathcal{N}}$ closer to the two-dimensional distribution than the total count comparisons, presumably because of the smaller number of counts per bin. For the $72$-mode, $216$-mode LS and $288$-mode experiments, this isn’t necessarily the case (see Appendix A), however all are closer to the thermalized ground truth than was observed in previous work [17] with threshold detector experiments.
Figure 5: Surface plots of two-dimensional GCP comparisons of theory and experiment for the $216$-mode HS experimental count patterns. Figures in the left column (a and c) plot of the full bivariate distribution $\mathcal{G}_{\{1,\dots,108\},\{109,\dots,216\}}^{(M)}(m_{1},m_{2})$, while the figures in the right column (b and d) plot the normalised difference $\Delta\mathcal{G}_{\{1,\dots,108\},\{109,\dots,216\}}^{(M)}(m_{1},m_{2})/\sigma$. In a) and b) the theoretical ideal ground truth is simulated, and in c) and d) the theoretical distribution is the thermalized ground truth. Thermalization component $\epsilon$ and measurement correction $t$, as well as the positive-P sample size are the same as in Fig.(3).
In these threshold detector experiments, $N_{E}\approx 4\rightarrow 5\times 10^{7}$ count patterns were generated and it was found that once the experimental data was binned in $d=4$ dimensions, there were too few counts per bin. This causes the bins to become more sparse, leading to a dramatic increase in experimental sampling errors that dominate statistical testing, outputting artificially small results. Therefore for these threshold detector experiments, reliable validation tests could only be performed for $d<4$ [17].
Almost all data sets from Borealis generated only $N_{E}\approx 1\times 10^{6}$ patterns. In this case, sampling errors dominate statistical testing at $d=2$, reducing the reliability of these tests such that only the $d=1$ comparisons are reliable.
The exception to this is the $16$-mode experiment whose $N_{E}\approx 8.4\times 10^{7}$ samples were used benchmark the system architecture through total variation distance tests. The network is small enough that count pattern probabilities can be computed up to a total of six detected photons. As the computational cost of computing probabilities grows with the total photon number, the full photon distribution was not computed in the analysis of Ref.[5].
Interestingly, the one-dimensional GCP for this data set is further from its ideal ground truth distribution than most experiments performed using the Borealis quantum network, where $\chi_{EI}^{2}/k\approx 2.5\times 10^{3}$ and $Z_{EI}=163$. Unlike data sets claiming quantum advantage, including decoherence and $\boldsymbol{T}$-matrix corrections is no longer enough to obtain agreement within sampling error between theoretical and experimental GCPs, instead producing statistical test results of $\chi_{ET}^{2}/k\approx 14\pm 2$ and $Z_{ET}\approx 18\pm 1$, although the difference here is mostly from the increased sample numbers.
@@ V.3 $d=2$ grouped count comparisons
An increase in the GCP dimension to $d=2$ shows the $16$-mode data is even further from both the ideal and non-ideal ground truth distributions as can be seen in Fig.(6), outputting $Z_{EI}=301$ and $Z_{ET}=86\pm 3$. Additional $d=2$ comparisons can be performed by randomly permuting the count patterns as each permutation allows one to compare different binned correlation moments. As some correlation moments may be closer to their theoretical values than others, Z-scores are likely to vary between permutations, as was observed in threshold detector experiment comparisons [17]. Focusing on comparisons with the non-ideal ground truth, a mean increase in the resulting Z-scores is observed from ten permutations, such that $\left\langle Z_{ET}\right\rangle_{\mathcal{RP}}\approx 139$, where $\left\langle\dots\right\rangle_{\mathcal{RP}}$ denotes a mean over random permutations. Although some permutations are closer to the non-ideal ground truth, with the closest being $Z_{ET}\approx 73$, the majority are further, where the largest observed Z-score was $Z_{ET}\approx 235$ from a single permutation.
@@
Although this is not unexpected, such results indicate further systematic errors besides those tested are present. These may include unequal losses in each network channel, phase noise, shot-to-shot fluctuations, miscounted photons, or detector efficiency fluctuations, among others. We conjecture that the presence of further systematic errors other than decoherence and measurement errors are also present in the $216$-mode HS and $288$-mode implementations. However the small sample size means such errors cannot be resolved in statistical testing for these data sets, limiting a thorough analysis.
Figure 6: Description follows Fig.(5) however in this case experimental count patterns are from the $16$-mode Borealis network, such that the $d=2$ GCP is $\mathcal{G}_{\{1,\dots,8\},\{9,\dots,16\}}^{(M)}(m_{1},m_{2})$. For this network, the thermalization component is $\epsilon=0.0648\pm 0.0005$ while the matrix measurement error is $t=0.9841\pm 0.0005$. All simulations used $E_{S}=1.2\times 10^{6}$ samples.
V.4 Photon number moments
@@
Unlike the GCP comparisons, decoherence and transmission matrix corrections no longer cause the theoretical moments to be within the sampling error-bars of the experimental moments, where $Z_{ET}\approx 45$ for the $216$-mode HS data set and the $288$-mode data set is $Z_{ET}\sigma_{\mathcal{N}}\approx 63\sigma_{\mathcal{N}}$ standard deviations away from its expected mean. Similar trends are present for all data sets where, as was the case for GCPs, the $16$-mode moments are furthest from their expected first order moments than any other experiment. We emphasize that statistical testing is more accurate for this data set due to the large number of experimental samples, allowing one to resolve finer differences between theoretical and experimental photon number moments.
@@
Figure 7: Upper panel: Input thermalised squeezed state theoretical mean output photon numbers $\left\langle\hat{n}^{\prime}_{j}\right\rangle$ versus mode number $j$ are compared to their experimental counterparts for data from the $216$-mode HS Borealis experiment, where simulations are performed for an ensemble size of $E_{S}=1.2\times 10^{6}$. Decoherence and $\boldsymbol{T}$-matrix parameters are $\epsilon=0.0510\pm 0.0005$ and $t=0.9941\pm 0.0005$ which, although improving the fit with experimental data, no longer cause theory and experiment to agree within sampling error. Lower panel: Corresponding normalised difference between marginal moments.
@@
Figure 8: Figure description follows Fig.(7), except data is now from the $288$-mode Borealis experiment where corresponding fitting parameters are $\epsilon=0.0547\pm 0.0005$ and $t=0.9848\pm 0.0005$.
@@ VI Conclusion
This method is computationally efficient, with computational run-times of a few minutes on a 12-core desktop at $\sim 10^{11}$ floating point operations/ sec (Linpack Flops) for one-dimensional GCP simulations of the largest sized networks implemented in the paper. Since the positive P-representation accurately computes correlation moments to any order generated in the photonic network, it is a time and resource efficient method of validation compared to classical algorithms that compute the Hafnian to replicate photo-detection measurements, or tensor network approximations [32].
Increasing the GCP dimension causes an increase in computation time, although this is still significantly faster than other validation methods. Directly calculating $100\times 100$ Hafnians are estimated [32] to take $14$ hours on Fugaku, one of the worlds largest supercomputers in 2024, at $\sim 10^{18}$ Flops. This would compute one probability, but rounding errors are unknown and are likely to be extremely large at current $64$ bit numerical precision. Generating enough classical counts for validation using this approach would take $\sim 10^{16}s$ [32] on Fugaku, indicating an overall speed-up of $>10^{18}$ using the positive-P method at these network sizes.
@@
Our method does not generate classical counts, so it is strictly for validation of the computational task, and does not replicate the required random counts. Validation tests were performed on GBS data from a recent experiment claiming quantum advantage [5]. We showed that within sampling error agreement between theory and experiment for one-dimensional, but not two-dimensional GCPs is obtained once decoherence is included in the input states, and a transmission matrix measurement error correction is applied. There are large discrepancies from $\chi^{2}$ tests on the individual channel count probabilities, because these can be calculated and measured very accurately. This result agrees with a trend obtained previously for threshold detector experiments [9, 17].
@@
Comparisons with threshold detector experiments are slightly misleading, however. Due to the smaller number of experimental samples from data sets claiming quantum advantage using PNR methods, only a small fraction of the possible tests could be performed. This was because experimental sampling errors were too large to apply more sensitive tests. Based on comparisons of the $16$-mode experimental data, whose sample size was large enough to perform accurate statistical testing, we conjecture that more experimental errors are present in the quantum advantage data sets than those tested for in this paper.
@@
In the longer term, due to its computational efficiency and ability to generate any measurable moment, the positive-P based binning method could be used as a means to provide error correction, either by correcting parameter values as shown in this paper, or by giving feedback so they can be adjusted. We suggest that this could provide a more practical route to quantum computational advantage than only using improved hardware, as is also thought to be the case for logic gate based quantum computers.
@@ Acknowledgments
The graphs presented in Section V of the main text compare theory and experiment predominantly for the $216$-mode HS and $288$-mode data sets, due to claims of quantum advantage in the original paper. Here we summarize statistical test results for all data sets in [5], starting with one-dimensional GCPs for the ideal ground truth in Table.(1). We use the subscripts $EI$ and $ET$ to denote comparisons between experimental data and the ideal and non-ideal ground truth distributions, respectively. These results clearly show that there are very large discrepancies in all data-sets, and that these differences cannot be explained by the intrinsic sampling error from using a finite number of samples.
A summary of the non-ideal ground truth comparisons are presented in Table (2), where the $t$ and $\epsilon$ fitting parameters are found using a Nelder-Mead simplex algorithm which minimizes the distance between experimental and theoretical distributions. As stated earlier, experimental data agrees with our modified theoretical predictions within sampling error for all data sets bar the $16$-mode experiment.
@@
Comparisons of the two-dimensional GCPs and mean output photon number moments are presented in Table.(Appendix A: Summary of all experimental comparisons) and Table.(4), respectively, for the ideal and non-ideal ground truth distributions. These show large differences in the two-dimensional GCPs compared to the ideal ground-truth. However, the two-parameter fitting model used in the one-dimensional comparisons is insufficient to give agreement in the thermalized case.
Mode number $Z_{EI}$ $k$ $\left|Z_{ET}\right|$ $k$
$16$ $301$ $426$ $86\pm 3$ $416$
          $72$ $82$ $772$ $12\pm 1$ $736$
          $216$-(LS) $75$ $658$ $10\pm 1$ $640$
          $216$-(HS) $58$ $4326$ $9\pm 1$ $4154$
          $288$ $203$ $4877$ $11\pm 1$ $4571$
Table 3: Summary of statistical tests for comparisons of two-dimensional GCPs $\mathcal{G}_{\{1,\dots,M/2\},\{M/2+1,\dots,M\}}^{(M)}(m_{1},m_{2})$ for all experimental data sets obtained from [5]. Chi-square and Z-statistic tests are generated from positive-P simulated GCPs with an ensemble size of $E_{S}=1.2\times 10^{6}$. The $\epsilon$ and $t$ fitting parameters are the same those used in Table.(2), where $Z_{ET}$ values are averages over ten simulation runs with corresponding standard deviations.
@@
The last table, giving the comparisons for the mean photon number per channel, has a much lower sampling error than the other results tabulated, since all the output samples can be used to obtain these mean values. This shows discrepancies outside the sampling error limits, indicating that transmission matrix parameter errors could be a significant part of the overall differences. This is consistent with the increased errors found for the two-dimensional GCPs.
Appendix B: Estimating sampling errors of moments
@@
Theoretical sampling errors are estimated numerically by subdividing the total phase-space ensemble as $E_{S}=N_{S}N_{R}$. Breaking the full ensemble of phase-space samples into sub-ensembles is a common numerical procedure for applications of phase-space representations yielding two complementary benefits: efficient multi-core parallel computing, decreasing computation time depending on available resources, and estimates of theoretical sampling errors [43].
The first sub-ensemble corresponds to sampling from the Gaussian phase-space distribution $N_{S}$ times and $N_{R}$ is the number of times the computation is repeated. Provided $N_{R}\gg 1$, the second sub-ensemble allows one to accurately estimate the standard deviation in the ensemble mean as $\sigma_{T,i}=\sigma_{t,i}/\sqrt{N_{R}}$ [52, 55], where $\sigma_{t}$ is the computed sample standard deviation which takes the general form [43]
@@
Figure 9: The exact mean photon number (solid black line) of $N=50$ thermal states sent into a lossless linear network compared to the mean photon number (solid orange dots) of a classical algorithm that replicates PNR measurements of classical states using the diagonal P-representation. The $N_{F}=4\times 10^{6}$ count patterns are binned to form the observable moments, and the standard deviation of these moments from their exact values $\sigma_{A}$ is the sampling error estimated using Eq.(45).